OpenStudy (anonymous):

A physical education class has three times as many girls as boys. During a class basketball game, the girls average 18 points each, and the class as a whole averages 17 points per person. How many points does each boy score on average?

OpenStudy (anonymous):

a trick is to use actual numbers. say there is only one boy and therefore there must be 3 girls, for a total of 4

OpenStudy (experimentx):

so quite opposite than the last illustration huh?

OpenStudy (anonymous):

then \[\frac{18\times 3+x}{4}=17\] \[54+x=68\] \[x=14\]

OpenStudy (anonymous):

so the lone boy scored 14 points. since it doesn't matter for the problem how many there are, this is the answer

OpenStudy (anonymous):

I got it now. Thank you

OpenStudy (ashtorah13):

but why do you multiply 18 and 3?